An instrument cabinet having a mass of 250 kg is r
Answer #1

hooke's law

F=-kx ...........................(1)

Second newton's law

F=m.a...........................(2)

simple harmonic motion

x=A sin(wt) ..................(3)

a= - A.w2sin(wt).............(4)

equaling (1) and (2)

-K.x=m.a .....................(5)

inserting (3) and (4) in (5)

-K [ A sin(wt) ]=m.[  A.w2sin(wt) ]

clearing for w

\omega=\sqrt{\frac{K}{m}}.............(6)

w=2. pi.f ..................(7)

equaling 6 and 7

2\pi.f=\sqrt{\frac{K}{m}}

clearing for f

f=\frac{1}{2\pi}\sqrt{\frac{K}{m}}

f=\frac{1}{2\pi}\sqrt{\frac{1000 000_{N/m}}{250_{kg}}}=10.065_{hz}

Yes, this cabinet has a natural oscillation frequency that places it at risk for resonant failure in an earthquake.

1) f > 12

m=250kg

(12_{hz}*2\pi)^2*250_{kg}=K>1421223N/m

2) f>12

k=1000000n/m

m=\frac{1000000_{N/m}}{(12_{hz}*2\pi)^2}<175.90_{kg}

If is possible decrease the mass of the cabinet below of 175.90kg it could be the solution, else increase the K spring over the 1422 KN/m